# Verification and Validation for Trustworthy Scientific Machine Learning

### v1

@@ 4.2 Solution verification
Code verification establishes a code’s ability to reproduce known idealized solutions, while solution verification, performed after code verification, assesses the code’s accuracy on more complex yet tractable problems defined by more realistic boundary conditions, forcing, and data. For example, code verification of ice sheets may use manufactured solutions, whereas solution verification may use more realistic MISMIP benchmarks [17]. In solution verification, the numerical solution cannot be compared to a known exact solution, and the convergence rate to a known solution cannot be established. Instead, solution verification must use other procedures to estimate the error introduced by the numerical discretization.
@@
While thorough solution verification of CSE models is challenging, these difficulties are further amplified for SciML models. Currently, solution verification of SciML models simply consists of evaluating a trained model’s performance using test data separate from the training data. However, this is insufficient as solution verification requires quantifying the impact of increasing data and model complexity on model error. Yet, unfortunately, performing a posteriori error estimation for many SciML models using techniques like Richardson extrapolation is difficult due to the confounding of model expressivity, statistical sampling errors, and variability introduced by converging to local solutions or saddle points of nonconvex optimization, making it challenging to monotonically decrease the error of SciML models such as neural networks.
Until convergence theory for SciML models improves and automated procedures are developed to change SciML model hyperparameters as data increases, solution verification of SciML models should repeat the sensitivity tests proposed for code verification (Section 4.1) with two key differences: First, verification experiments used to generate verification data must be specifically designed for solution verification, as not all verification data equally informs solution verification efforts, similar to observations made when creating validation datasets for CSE models [72]. See Section 6.1 for more information on important properties of verification benchmarks. Second, while ideally the convergence of SciML errors on realistic benchmarks should be investigated, it may be computationally impractical. Thus, solution verification should prioritize quantifying errors using the model complexity and data amount that will be used when deploying the SciML model to its application domain.
@@ 5.1 Calibration
Once a CSE model has been verified, it must be calibrated to match experimental data that contains observational noise. This calibration requires solving an inverse problem [100], which can be either deterministic or statistical (e.g., Bayesian). The deterministic approach formulates the inverse problem as a (nonlinear) optimization problem that minimizes the mismatch between model and experimental data. This formulation requires regularization to ensure well-posedness, typically chosen using the L-curve [35] or the Morozov discrepancy principle [3]. The Bayesian approach replaces the misfit with a likelihood function based on the noise model, while using a prior distribution for regularization. This prior distribution ensures well-posedness while encoding typical parameter ranges and correlation lengths. We recommend Bayesian methods for calibration because they provide insight into the uncertainty of the reconstructed model parameters.
The calibration of SciML operator, reduced-order, and hybrid CSE-SciML models is distinct from SciML training and follows similar principles to CSE model calibration. These models are first trained using simulation data for solution verification. Next, observational data (called calibration data) determines the optimal model input values that match experimental outputs. For instance, calibrating a SciML ice-sheet model such as that of Ref. [37] requires finding optimal friction field parameters of a trained SciML model, which best predict observed glacier surface velocities, given the noise in the observational data.
Calibration typically improves a model’s predictive accuracy on its application domain, but the informative value of calibration data varies significantly. Therefore, researchers should select calibration data strategically to maximize information content within their experimental budget. See Section 6.1 for further discussion on collecting informative data.
@@ 5.2 Model validation
Model validation is the “substantiation that a model within its domain of applicability possesses a satisfactory range of accuracy consistent with the intended application of the model” [83]. Validation involves comparing computational results with observational data, then determining if the agreement meets the model’s intended purpose [56]. For CSE models with unacceptable validation agreement, modelers must either collect additional calibration data or refine the model structure until reaching acceptable accuracy. SciML models follow a similar iterative process but offer an additional option: to collect more training data.
Model validation must occur after calibration and requires independent data not used for calibration or training. For our conceptual ice-sheet model, calibration matches surface velocities assumed to represent pre-industrial conditions, while validation assesses the calibrated model’s ability to predict grounding line change rates at the start of this decade. Performance metrics must target the specific modeling purpose. For optimization tasks, metrics should measure the distance from true optima obtained via the SciML model or bound the associated error [14]. For uncertainty estimation, metrics should quantify errors in uncertainty statistics through moment discrepancies or density-based measures like (shifted) reverse and forward Kullback–Leibler divergences. For explanatory SciML modeling, validation metrics must also assess physical fidelity: adherence to physical laws, conservation properties (such as mass and energy), and other constraints. As with verification, the validation concept should encompass data validation, particularly whether training data adequately represents the application space.
Validation determines whether a model is acceptable for its specific purpose rather than universally correct. The definition of acceptable is subjective, depending on validation metrics and accuracy requirements established by model stakeholders in alignment with the problem definition and model purpose (see Section 3). Moreoever, validation itself does not constitute final model acceptance, which must be based on model accuracy in the application domain, as discussed in Section 5.3.
Two additional considerations complete our discussion of model validation. First, this validation differs from the concept of cross validation, which estimates ML model accuracy on data representative of the training domain during development. The validation described here assesses accuracy in a separate validation domain. Second, validation data varies in informative value. Validation experiments should “capture the essential physics of interest, including all relevant physical modeling data and initial and boundary conditions required by the code” [73]. Most critically, validation data must remain independent from training and calibration data.
@@ 5.3 Prediction
SciML models, like CSE models, are subject to numerous sources of uncertainty. The impact of these uncertainties on model predictions must be quantified. Several sources of uncertainty affect CSE models. These include: numerical errors, from approximating the solution to governing equations; input uncertainty arises, which is caused by inexact knowledge of model inputs; parameter uncertainty, which stems from inexact knowledge of model coefficients; and model structure error representing the difference between the model and reality. SciML models contain all these uncertainties. They also incorporate additional uncertainties from sampling and optimization errors, as discussed previously.
@@
Sampling error arises from training a model with a finite amount of possibly noisy data. For a fixed ML model structure and zero optimization error, this error decreases as the amount of data increases. Optimization error represents the difference between the optimized solution, which is often a local optimum, and the global solution for fixed training data. Optimization error can enter CSE models during calibration. Optimization error affects SciML models more significantly because it occurs both during calibration and training. Linear approximations, for example, based on polynomials, achieve zero optimization error during training to machine precision. However, nonlinear approximations such as neural networks often produce non-trivial optimization errors. Stochastic gradient descent demonstrates this by producing different parameter estimates due to stochastic optimization randomness and initial guesses.
The identified sources of modeling uncertainty require parameterization for sampling. Expert knowledge typically guides the construction of prior distributions that represent parametric uncertainty. This parameterization should occur during problem definition. Bayesian calibration updates these priors into posterior distributions using calibration data. The model must then propagate all uncertainties onto predictions in the application domain. Monte Carlo quadrature accomplishes this propagation by drawing random samples from the uncertainty distributions. The method collects model predictions at these samples and computes empirical estimates of important statistics defined by validation criteria, such as mean and variance.
We emphasize the impact of all sources of error and uncertainty must be quantified. Simply estimating the impact of error caused by using finite sample sets, for example estimated by generative models such as variational autoencoders of Gaussian processes is insufficient. Moreover, complete elimination of uncertainty is impossible. Consequently, model acceptance, like validation, must rely on subjective accuracy criteria established through stakeholder communication. For instance, acceptance criteria for predicted sea-level change from melting ice-sheets at year 2100 may specify that prediction precision reaches 1% of the mean value. Yet, engineering applications, such as those focused on aerospace design, may have much higher accuracy requirements.
The aforementioned Monte-Carlo based UQ procedure effectively quantifies the impact of parameterized uncertainties on model predictions. However, model structure error remains difficult to parameterize in both SciML and CSE modeling. Validation can partially assess model structure error. However, experiments rarely cover all conditions of use. Specifically, validation tests only the model’s interpolation ability within the convex hull of available data and assumptions. This limitation creates challenges when applying the model outside its validation domain. Some progress exists in quantifying extrapolation error for “models based upon highly-reliable theory that is augmented with less-reliable embedded models” [77]. However, such hybrid CSE-SciML models rely on well-established physics-based governing equations to support extrapolation confidence. Pure SciML models still require substantial research to develop reliable methods for estimating model structure uncertainty.
@@ 6.1 Data Sources
In addition to listing prior knowledge that can be used to build an ML model, the types of data used to train a model affects model V&V. The data sources, which will be used to train the SciML model, must be carefully chosen and the reasons for those choices must be documented. Important data properties to consider are data fidelity and bias, data noise and the cost of collecting training data, which limits the amount that can be collected. Some studies make recommendations on the amount of data that should be used to train an ML model based on the number of model input features [117]. However, we believe the amount of data should be defined by the accuracy requirements of the intended use of the SciML model. Future work on generalization accuracy and statistical learning theory for SciML models will help better elucidate these mathematical trade-offs.
When training a SciML model, the fidelity of data impacts the performance of a model. For example, when training using simulation data, the numerical discretization of the model generating the data is important. Ideally, training data should be generated using the most accurate (highest-fidelity) discretization. However, often a cheaper but less accurate discretization (lower-fidelity) is used because it allows for more training data to be collected. While more low-fidelity training data allows a better approximation of the low-fidelity simulation model, it introduces a bias in the SciML model relative to the high-fidelity model. The impact of this error on quantities predicted by a model, for instance the optimal objective of a design problem, must be quantified. Alternatively, multi-fidelity SciML models, which combine limited high-fidelity data with larger amounts of low-fidelity data [29, 67, 19], can sometimes be used to mitigate this issue.
The creation of informative benchmarks can substantially improve the quality of V&V analyses and enable easier comparison of methods and SciML structures. However, unlike traditional ML, SciML, and CSE, lack a set of commonly-used and agreed-upon benchmark problems, for both solution verification and validation. Consequently, further work is needed to complement early efforts to create SciML benchmarks [102]. Motivated by the work of Oberkampf [74], we believe the purpose and details of the benchmark should be precisely stated, including how comparisons with the benchmark must be made and the the exact verification or validation acceptance criteria. For example, a natural acceptance criteria for a SciML surrogate used for UQ is that the surrogate can accurately compute the mean and variance of the true model it replaces to a specified level of accuracy. In contrast, a reasonable acceptance criterion of a SciML model used for optimization would be based on the comparison of the optimal solution obtained by the SciML model with known or trusted reference optima. Lastly, benchmarks must be made widely accessible, though this can be difficult when some benchmarks are associated with terabytes of data. Thus, until community benchmarks and mechanisms for their dissemination are developed, it is crucial that researchers thoroughly document these choices and explain the significance of the problems they create.
@@ 6.5 Software quality assurance and reproducibility
SQA requires both static analysis [62, 25] and dynamic testing [48]. Static analysis checks for errors without executing code and can be applied to an entire code base before dynamic tests are executed. This is useful because runtime tools, such as Valgrind [71], only test the code that is executed and thus can miss corner cases that are not tested [32]. Dynamic testing comprises unit tests, integration tests and system tests. All of these software engineering practices have low adoption in ML software [97]. Software engineering is particularly challenging in SciML models because it has “…all of the problems of non-ML software systems plus an additional set of ML-specific issues” [108]. While deterministic CSE can often use theoretically derived test “oracles” that capture the expected behavior of a model [49], ML’s stochastic nature requires aggregating multiple model runs to assess variability. System-wide testing is often needed to understand emergent properties, making unit testing less useful. Additionally, code coverage metrics are less meaningful for SciML since decision logic is relearned with each training [116]. Despite the importance of SQA, it is often neglected in both CSE and SciML. An analysis of 62 scientific software studies found that most used only one type of testing, and none used all types (unit, system, integration, acceptance, and regression testing) [50]. Improving SQA adoption is thus crucial for building trustworthy SciML models. SciML studies should make every effort to establish confidence in the software implementation, via modular tests and benchmarks, and adoption of software engineering best practices.
Reproducibility is essential for credibility building of computational models [70], and can be understood as another form of quality assurance. For reproducibility, authors should provide comprehensive documentation, including data specifications, computing requirements, software versions, and public code repositories. The optimized parameter and hyper-parameters of SciML models, e.g., the width and depth of a neural network and its weights and biases, should also be released. Repositories should be well-organized with tutorials and reproduction scripts [51], and standard public licenses should be adopted that stipulate reuse rights of the artifacts.
Warnings about lack of reproducibility in ML research have been raised before [43, 33]. A common failing, with an obvious solution, is not sharing the source code associated with the published research. Yet simply depositing the software in a public repository is far from sufficient to ensure reproducibility in ML research. Rigorous documentation of the training pipeline is necessary, including details of the optimization and regularization applied, activation functions, hyperparameters, learning-rate schedules, and more. Data provenance, as discussed earlier, should be managed strictly and documented; ideally the full data sets should be deposited in trusted repositories that provide a persistent identifier. None of these efforts is minor, and sometimes they may seem unrewarding, yet they are necessary, and increasingly paid attention to by funders, journals, and conferences. Higher reproducibility standards will advance SciML’s trustworthiness, and benefit authors by reduced barriers to the adoption of their proposed methods.
@@ 6.6 Cost-accuracy comparison with alternatives
SciML models must be compared with state-of-the-art alternatives throughout model development, including verification, validation, and prediction stages. However, a review of environmental SciML studies found that 37.8% made no ML comparisons, 44.6% compared with one SciML method, and only 14.9% compared with multiple methods [117].
@@
Although accuracy is crucial in method comparison, computational costs must also be documented, yet total modeling costs are often underreported. Researchers should detail all costs, including data generation (e.g., running a CSE model), parameter optimization, and prediction. They should also compare the total of these costs against state-of-the-art alternatives. For instance, in PDE-constrained optimization, direct PDE optimization costs should be compared to combined surrogate training and optimization costs.
Dismissing training costs as “offline” is misleading, however, these costs can be amortized when SciML models are used across multiple studies, as in learning to optimize [15]. Nevertheless, such amortization claims need qualification since SciML models’ cost-effectiveness—the trade-off between accuracy, development time, and cost—is harder to predict than for CSE models [96].
Finally, every effort must be made to avoid “the misuse of language, intentionally or unintentionally, as this affects the interpretation of study findings and has been described as “spin” [2] or reporting bias. While this may appear obvious, a recent review of 152 studies of clinical prediction using ML models found that 95% of abstracts that recommended the use of an ML in daily practice lacked any external validation of the developed models. Moreover, only 55.6% of studies recommended the clinical use of an ML model in their main text without any external validation. McGreivy and Hakim [66] found that 79% of 76 ML papers solving fluid PDEs used weak baselines when claiming to outperform standard methods. The same study also found that outcome reporting bias and publication bias are widespread. Of 232 abstracts surveyed, 94.8% reported only positive results, 5.2% reported mixed results, and none reported only negative results, suggesting reporting bias.

### v2

@@ 4.2 Solution verification
Code verification establishes a code’s ability to reproduce known idealized solutions, while solution verification, performed after code verification, assesses the code’s accuracy on more complex yet tractable problems defined by more realistic boundary conditions, forcing, and data. For example, ice sheets code verification may use manufactured solutions, whereas solution verification may use more realistic MISMIP benchmarks [14]. In solution verification, the numerical solution cannot be compared to a known exact solution and convergence rate, thus other procedures must be used to estimate the error introduced by numerical discretization.
@@
While thorough solution verification of CSE models is challenging, it is even more difficult for SciML models. Currently, solution verification of SciML models simply consists of evaluating a trained model’s performance on test data separate from the training data, which is insufficient as it does not quantify the impact of increasing data and model complexity on model error. A posteriori error estimation using techniques like Richardson extrapolation is complicated by model expressivity, statistical sampling errors, and variability from converging to local solutions in nonconvex optimization.
Until convergence theory for SciML models improves and automated procedures for adjusting hyperparameters as data increases are developed, solution verification should repeat the sensitivity tests proposed for code verification (Section 4.1), with two key differences: First, verification experiments must be specifically designed for solution verification, as not all verification data equally informs solution verification efforts, similar to observations made when creating validation datasets for CSE models [59] and discussed further in Section 6.1. Second, while investigating the convergence of SciML errors on realistic benchmarks is ideal, it may be computationally impractical; thus, solution verification should prioritize quantifying errors based on the model complexity and data amount intended for deployment.
@@ 5.1 Calibration
Once a CSE model has been verified, it must be calibrated to match experimental data that contains observational noise, which involves solving either a deterministic or statistical (e.g., Bayesian) inverse problem [87]. The deterministic approach formulates this as a (nonlinear) optimization problem that minimizes the mismatch between model and experimental data, requiring regularization methods like the L-curve [26] or the Morozov discrepancy principle [3]. In contrast, the Bayesian approach replaces the misfit with a likelihood function, based on the noise model, while using a prior distribution for regularization. We recommend Bayesian methods for calibration because they provide insight into the uncertainty of the reconstructed model parameters.
Calibrating SciML and hybrid CSE-SciML models is distinct from SciML training and follows similar principles to CSE model calibration. These models are first trained using simulation data for solution verification. Next, observational data (called calibration data) determines the optimal model input values that match experimental outputs. For instance, calibrating a SciML ice-sheet model, e.g. [28], requires finding optimal friction field parameters of a trained SciML model, which best predict observed glacier surface velocities, accounting for observational noise.
While, calibration typically improves a model’s predictive accuracy on its application domain, the informative value of calibration data varies significantly. Therefore, researchers should select calibration data strategically to maximize information content within their experimental budget. See Section 6.1 for further discussion on collecting informative data.
@@ 5.2 Model validation
Model validation is the “substantiation that a model within its domain of applicability possesses a satisfactory range of accuracy consistent with the intended application of the model” [70]. It involves comparing computational results with independent observational data, to determine if the agreement meets the model’s intended purpose [45]. For CSE models with unacceptable validation agreement, modelers must either collect additional calibration data or refine the model structure until reaching acceptable accuracy, while SciML modelers can also collect more training data.
Validation must occur after calibration and requires independent data not used for calibration or training and well chosen validation criteria [53]. In our conceptual ice-sheet model, calibration matches surface velocities representing pre-industrial conditions, while validation assesses the calibrated model’s ability to predict grounding line change rates at the start of this decade. Performance metrics must target the specific modeling purposes, such as measuring distance from true optima for optimization tasks [10] or quantifying errors in uncertainty statistics. For explanatory SciML modeling, validation metrics must also assess physical fidelity, including adherence to physical laws and conservation properties. As with verification, the validation concept should encompass data validation, particularly whether training data adequately represents the application space.
Note that the term validation, as used here, differs from the concept of cross validation, which estimates ML model accuracy on data representative of the training domain during development. Validation determines whether a model is acceptable for its specific purpose rather than universally correct. The definition of acceptable is subjective, depending on validation metrics and accuracy requirements established by model stakeholders in alignment with the problem definition and model purpose (see Section 3). The amount of data needed to meet these criteria will depend on the information content of the calibration data, which can vary substantially if not chosen judiciously. Validation experiments should “capture the essential physics of interest, including all relevant physical modeling data and initial and boundary conditions required by the code” [60]. Lastly, validation itself does not constitute final model acceptance, which must be based on model accuracy in the application domain, as discussed in Section 5.3.
@@ 5.3 Prediction
SciML models, like CSE models, are subject to numerous sources of uncertainty that must be quantified, including: numerical errors, which arise from approximating the solution to governing equations; input uncertainty, which is caused by inexact knowledge of model inputs; parameter uncertainty, which stems from inexact knowledge of model coefficients; and model structure error representing the difference between the model and reality. However, SciML models are also subject to sampling and optimization errors; sampling error arises from training with finite, potentially noisy data, while optimization error reflects the difference between local and global solutions, affecting SciML models during both calibration and training. Linear approximations, for example, based on polynomials, achieve zero optimization error during training to machine precision. However, nonlinear approximations such as neural networks often produce non-trivial optimization errors. Stochastic gradient descent demonstrates this by producing different parameter estimates due to stochastic optimization randomness and initial guesses.
All uncertainties must be propagated onto predictions made in the application domain. Quantifying a single error, such as those reported by generative models like variational autoencoders or Gaussian processes, is typically insufficient. All sources of uncertainty should be parameterized using expert knowledge to construct prior distributions, which are then updated through Bayesian inference and calibration data to form posterior distributions. Methods like Monte Carlo quadrature can randomly sample from these posterior distributions to compute empirical estimates of important statistics defined by validation criteria, such as mean and variance.
While Monte Carlo-based UQ procedures effectively quantify parameterized uncertainties, model structure error remains challenging to parameterize, and validation can only partially assess it, as experiments rarely cover all conditions of use. Progress has been made in quantifying extrapolation error for hybrid models that combine reliable physics-based equations with less-reliable embedded models [63], but pure SciML models still require significant research to develop reliable methods for estimating model structure uncertainty. Moreover, complete elimination of uncertainty is impossible. Consequently, model acceptance, like validation, must rely on subjective accuracy criteria established through stakeholder communication. For example, acceptance criteria for predicted sea-level change from melting ice sheets by 2100 may specify a prediction precision of 1% of the mean value, while engineering applications, such as aerospace design, may demand much higher accuracy requirements.
@@ 6.1 Data Sources
The types of data used to train a SciML model significantly affect V&V, so data sources must be carefully selected and documented. Key properties to consider include data fidelity, bias, noise, and the cost of data collection, which can limit the amount available. While some studies recommend data quantity based on the number of input features [104], we believe it should be determined by the accuracy requirements of the model’s intended use. Future work on generalization accuracy and statistical learning theory for SciML models will help better elucidate these mathematical trade-offs. Additionally, the fidelity of training data impacts model performance; for instance, using lower-fidelity simulation data may allow for more training samples but introduces bias relative to high-fidelity models. This error’s impact on predicted outcomes, such as the final objective value of an optimal design, must be quantified. Alternatively, multi-fidelity SciML models [16] can mitigate this issue by combining limited high-fidelity data with larger amounts of low-fidelity data.
Creating informative benchmarks can significantly enhance the quality of V&V analyses and facilitate comparisons of methods and SciML structures. However, unlike traditional ML, SciML and CSE lack commonly agreed-upon benchmark problems for solution verification and validation, necessitating further work to build on initial efforts to establish SciML benchmarks [89]. Following Oberkampf’s guidance [61], benchmarks should clearly state their purpose, comparison methods, and specific acceptance criteria; for instance, a SciML surrogate for uncertainty quantification should accurately compute the mean and variance of the true model to a specified level of accuracy, while a SciML model for optimization should compare its optimal solution to trusted reference optima. Additionally, benchmarks must be widely accessible, though this can be challenging with large datasets. Until community benchmarks and dissemination mechanisms are developed, researchers should thoroughly document their choices and explain the significance of the problems they address.
@@ 6.5 Software quality assurance and reproducibility
SQA requires both static analysis [50] and dynamic testing [37]. Static analysis checks for errors without executing code and can be applied to an entire code base before dynamic tests are executed. This is useful because runtime tools only test the code that is executed and thus can miss corner cases that are not tested [23]. Dynamic testing comprises unit tests, integration tests and system tests. Despite the importance of SQA, all of these software engineering practices have low adoption in ML [84] and CSE software [40]. Software engineering is particularly challenging in SciML models because it has “…all of the problems of non-ML software systems plus an additional set of ML-specific issues” [94]. While deterministic CSE can often use theoretically derived test “oracles” that capture the expected behavior of a model [39], ML’s stochastic nature requires aggregating multiple model runs to assess variability. System-wide testing is often needed to understand emergent properties, making unit testing less useful. Additionally, code coverage metrics are less meaningful for SciML since decision logic is relearned with each training [103]. Improving SQA adoption is thus crucial for building trustworthy SciML models. SciML studies should make every effort to establish confidence in the software implementation, via modular tests and benchmarks, and adoption of software engineering best practices.
Reproducibility is essential for building the credibility of computational models [58] and serves as a form of quality assurance, which is often missing from ML research [32, 24]. Rigorous documentation of the training pipeline is necessary, including details of the optimization and regularization applied, activation functions, hyperparameters, learning-rate schedules, data specifications, computing requirements, software versions, and public code repositories. Repositories must be well-organized with tutorials and reproduction scripts [41], store the optimized parameters and hyperparameters of SciML models, such as neural network architecture and weights, and standard public licenses should state reuse rights. Simply depositing source code in a public repository is insufficient. Data provenance must be strictly managed, with full datasets ideally deposited in trusted repositories that offer persistent identifiers. While these efforts may seem burdensome, they are crucial for enhancing SciML’s trustworthiness and are increasingly emphasized by funders, journals, and conferences, ultimately benefiting authors by lowering barriers to the adoption of their methods.
@@ 6.6 Cost-accuracy comparison with alternatives
SciML models must be compared with state-of-the-art alternatives throughout model development, including verification, validation, and prediction stages. However, such comparisons are often neglected in literature [78]. For example, a review of environmental SciML studies found that 37.8% made no ML comparisons, 44.6% compared with one SciML method, and only 14.9% compared with multiple methods [104].
@@
Although accuracy is crucial in method comparison, computational costs must also be documented, yet total modeling costs are often underreported. Researchers should detail all costs, including data generation (e.g., running a CSE model), parameter optimization, and prediction. For instance, in PDE-constrained optimization, direct PDE optimization costs should be compared to combined surrogate training and optimization costs. Additionally, dismissing training costs as “offline” is misleading, however, these costs can be amortized when SciML models are used across multiple studies, as in learning to optimize [12]. Nevertheless, such amortization claims need qualification since SciML models’ cost-effectiveness—the trade-off between accuracy, development time, and cost—is harder to predict than for CSE models [83].
Finally, efforts must be made to avoid the misuse of language, as it can distort the interpretation of study findings and lead to ”spin” or reporting bias [2]. For example, McGreivy and Hakim [55] found that 79% of 76 ML papers on fluid PDEs used weak baselines to claim superiority over standard methods, and of 232 abstracts surveyed 94.8% reported only positive results, indicating widespread reporting bias.